OpenStudy (anonymous):

Please help!! fans/medals Jennifer has been saving for college for 57 months. The first month she saved $11. She was able to save more money each month than the month before. She ended up saving $19,779.00. How much more did she save each month?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

" She was able to save more money each month than the month before." Assume her monthly savings follows an arithmetic sequence and that she saves "d" dollars more each month: 11, 11+d, 11+2d, 11+3d, ......., 11+56d The last term is 11+56d because she saves for 57 months. The sum of an arithmetic progression is: (first term + last term) / 2 * n where n is the total number of terms which is 57 Sum = { 11 + 11 + 56d } / 2 * 57 = 19,779 (11 + 28d) * 57 = 19779 divide both sides by 57. The subtract 11 from both sides divide both sides by 28 and solve for d.